# Hanke-Raus heuristic rule for iteratively regularized stochastic gradient descent

### v1

@@
Abstract. In this work, we present a novel variant of the stochastic gradient descent method termed as iteratively regularized stochastic gradient descent (IRSGD) method to solve nonlinear ill-posed problems in Hilbert spaces. Under standard assumptions, we demonstrate that the mean square iteration error of the method converges to zero for exact data. In the presence of noisy data, we first propose a heuristic parameter choice rule (HPCR) based on the method suggested by Hanke and Raus, and then apply the IRSGD method in combination with HPCR. Precisely, HPCR selects the regularization parameter without requiring any a-priori knowledge of the noise level. We show that the method terminates in finitely many steps in case of noisy data and has regularizing features. Further, we discuss the convergence rates of the method using well-known source and other related conditions under HPCR as well as discrepancy principle. To the best of our knowledge, this is the first work that establishes both the regularization properties and convergence rates of a stochastic gradient method using a heuristic type rule in the setting of infinite-dimensional Hilbert spaces. Finally, we provide the numerical experiments to showcase the practical efficacy of the proposed method.
Keywords: Stochastic gradient descent; Iterative Landweber method; Convergence rates; Variational inequalities; Inverse problems; Nonlinear ill-posed problems; Regularization.
@@ 1. Introduction
This work focuses on deriving an approximate solution for a system of nonlinear ill-posed equations of the form
((1.1)) $F_{i}(u)=y^{\dagger}_{i},\hskip 5.69054pti=0,1,\ldots,P-1,$
[…]
@@
where $F:U\rightarrow Y^{P}$ is defined as
$F(u)=\begin{pmatrix}F_{0}(u)\\ \vdots\\ F_{P-1}(u)\end{pmatrix}\ \ \text{and}\ \hskip 5.69054pty^{\dagger}=\begin{pmatrix}y^{\dagger}_{0}\\ \vdots\\ y^{\dagger}_{P-1}\end{pmatrix}.$
@@
Usually, exact data $y^{\dagger}$ is never precisely available. Instead, a noisy data $y^{\delta}\in Y^{P}$ is available such that
((1.3)) $\|y^{\dagger}-y^{\delta}\|\leq\delta,$
for noise level $\delta\geq 0$. Here, we write $\delta=\sqrt{\delta_{0}^{2}+\delta_{1}^{2}+\cdots+\delta_{P-1}^{2}}$ and $y^{\delta}=(y^{\delta}_{0},y^{\delta}_{1},...,y^{\delta}_{P-1})$, where
((1.4)) $\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\delta_{i}\ \ \text{for}\ \ i=0,1,2,\cdots,P-1.$
@@
Due to ill-posedness of (1.1), its solution may not exist and even if it exists, it may not be unique. Moreover, the solution(s) can be highly sensitive to perturbations in the noisy data $y^{\delta}$ (see [6]). Consequently, regularization methods are needed for deducing the stable approximate solution of (1.1).
@@
One of the prominent and highly effective classes of regularization methods is that of iterative regularization. This approach has achieved significant success in solving a wide variety of inverse problems (see [6, 16] and the references therein). Among these methods, the Landweber Iteration Method (LIM) is a well-studied and widely recognized classical iterative technique. For a Fréchet differentiable (or smooth) forward mapping $F:U\rightarrow Y$, LIM can be written as
((1.5)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta}),\hskip 28.45274ptk\geq 0.$
@@
Here $u_{0}^{\delta}:=u_{0}$represents an initial guess that incorporates the prior knowledge about the solution to be reconstructed and $F^{\prime}(u_{k}^{\delta})^{*}$ is the adjoint of the Fréchet derivative of $F$ at $u_{k}^{\delta}$.
@@
Scherzer [26] introduced a modification to the method (1.5) by introducing a damping term and named it as iteratively regularized Landweber iteration method. This method can be written as
((1.6)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta})-\lambda_{k}(u_{k}^{\delta}-u^{(0)}),\hskip 28.45274ptk\geq 0,$
where $\lambda_{k}\geq 0$. In comparison to the method (1.5), the convergence rates analysis of the method (1.6) needs fewer assumptions on the mapping $F$. Furthermore, the method (1.6) converges to a solution which is closed to $u^{(0)}$, however, the method (1.5) needs additional assumptions for the similar convergence behaviour (see [16] for nice details). This means that the inclusion of the additional damping term in (1.6) has several advantages over the method (1.5). When the noisy data is available, a proper stopping rule (a-posteriori) must be employed in order to show that the iterative schemes (1.5) and (1.6) are a regularization one. The commonly incorporated stopping rule is well known as discrepancy principle, i.e., the method is halted after $k_{\delta}=k_{\delta}(\delta,y^{\delta})$ steps, where
@@
for some $\tau>1.$ In real-world applications, noise level is not always available or trustworthy, which makes the discrepancy principle (1.7) difficult to use. Consequently, while applying the discrepancy principle, overestimation or underestimation of the noise level may result in significant degradation of the reconstruction accuracy. Therefore, it is reasonable to work on heuristic stopping rules that do not rely on the knowledge of noise level. Real [25] proposed a heuristic rule for the Landweber iteration (1.5) in Banach space setting based on the discrepancy principle, by taking motivation from the method of Hanke and Raus [9]. For (1.5), heuristic rule determines an integer $k_{*}:=k_{*}(y^{\delta})$ by minimizing
@@
over $k$, where $a\geq 1$ is a fixed number and $p>0$ is a constant.
It can be observed that for the conventional iterative regularization methods for solving inverse problems (such as the Landweber method [16], Levenberg-Marquardt method [5], conjugate gradient method [16], non-stationary Tikhonov iterative regularization method [20] and iteratively regularized Gauss-Newton method [16], among others), one of the common challenges is the significant computational burden at each iteration, mainly due to the need to process all the data, which can be excessively large and resource-intensive. To address this issue, a potential and effective approach is the utilization of stochastic gradient descent (SGD), a method introduced by Robbins et al. [24]. Further, Jin et al. [13] studied the convergence analysis of SGD method and demonstrated its regularizing nature when there is a noise in the data. The basic version of SGD can be written as (see [14])
@@
where $\mu_{k}$ is the corresponding step size and $i_{k}$ is the uniformly drawn index from the index set $\{0,1,...,P-1\}.$ It can be noted that the method (1.9) is a randomized version of (1.5). For (1.9), the stopping index $k(\delta)\in\mathbb{N}$ is defined using the following a-priori stopping rule
@@
We refer to [14, 17] for some recent literature on SGD method using a-priori stopping rule for linear as well nonlinear inverse problems and [12, 27] on SGD method using a-posteriori stopping rule for linear inverse problems. It is evident from these references that they exclusively use an a-priori or a-posteriori stopping rule in developing their convergence theories. From this observation, it is of significant interest to develop heuristic rules for stochastic gradient descent (SGD) type methods. Therefore, the primary objective of this paper is to explore the convergence analysis of the following iteratively regularized stochastic gradient descent (IRSGD) under heuristic stopping rules:
@@
where $\omega_{k}$ is the step size, $\lambda_{k}$ is the weighted parameter, $F^{\prime}_{i}(u)$ is the Fréchet derivative of $F_{i}(u)$, $i_{k}$ is an index randomly drawn from the index set $\{0,1,\ldots,P-1\}$ through uniform distributation. This method differs from the SGD method in the sense that an extra damping term $-\lambda_{k}(u_{k}^{\delta}-u_{0})$ is considered in its formulation. A key advantage of studying IRSGD lies in its ability to reflect SGD by setting the damping factor $\lambda_{k}=0$. Our approach is novel in the realm of stochastic methods for solving ill-posed inverse problems as it provides convergence analysis without relying on an a-priori stopping rule. Additionally, it circumvents the need to know the noise level, which further enhances its practical applicability.
For SGD method, Jin et al. [14] established the convergence rate in the mean squared norm in the context of nonlinear ill-posed problems, assuming an appropriate source condition. Further, Jahn et al. [12] further extended these results using an a-posteriori stopping rule, however, precise estimates for the convergence rates are not provided in that work. Inspired by these findings, one of the main aims of this work is to develop convergence rates for the iterates generated by (1.11) under an a-posteriori stopping rule. In addition to that, we also derive convergence rates under heuristic stopping criteria. To the best of our knowledge, this is the first study that addresses convergence rates of a stochastic type method for solving nonlinear ill-posed problems under heuristic stopping rule.
The main contributions of this paper are outlined as follows:
  • Instead of focusing on the analysis of a single equation, we turn our attention to exploring stochastic methods for solving systems comprising a large number of equations (i.e., $P$ is large in (1.1)). Stochastic methods involve the random consideration of each equation within equation (1.1) independently, thereby significantly reducing memory requirements. This approach finds relevance in numerous real-world scenarios (see [7, 21, 22]).
  • The method is fully data-driven, as it is applicable without using a-priori or a-posteriori stopping rule. This is the first attempt of construction of a stochastic method using heuristic type stopping rule. We establish the (strong) convergence of IRSGD method, which improves the weak convergence (i.e., convergence in probability) results provided in [12].
  • We establish the convergence rates for the proposed scheme (1.11) using a step size $\omega_{k}\in[\omega,\Omega]$ and a decaying weight parameter, under two types of stopping criteria: the discrepancy principle and a heuristic rule. The rates are comparable to those provided in [14, Theorem 2.5], [8, Theorem 3.2], and [13, Theorem 2.2] (see Table 1).
  • This work presents experimental evidence showcasing the method’s enhanced efficiency and accelerated performance achieved through the inclusion of a damping factor in SGD. This modification yields significantly better results compared to the standard SGD approach discussed in [14].
The rest of the paper is structured as follows. In Section 2, we list the preliminaries needed in our work. Section 3 covers the convergence analysis of our method in the presence of noisy and exact data. Section 4 discusses the convergence rates of our method for a-posteriori and heuristic stopping rules under certain assumptions. Section 5 presents numerical experiments on a system of ill-posed problems, demonstrating the applicability of the proposed scheme with heuristic stopping rule. The final section offers concluding remarks. Additionally, the Appendix includes few auxiliary lemmas required for our analysis.
2. Preliminaries
In this section, we review certain basic definitions and notation related to our work, see [2, 4, 14, 15, 23] for more details.
The notation $u_{k}$ is used to denote the iterates of our method corresponding to the exact data $y^{\dagger}.$ Additionally, we refer $\mathcal{F}_{k}$ as the filtration generated by the random indices $\{i_{1},...,i_{k-1}\}$ up to the $(k-1)$th iteration. Because we randomly select the indices $i_{k}$, the iteration $u_{k}^{\delta}$ in the IRSGD method becomes a random process. To assess its convergence, multiple methods are available. In our case, we will use the mean squared norm defined by $\mathbb{E}[\|\cdot\|^{2}]$, where $\mathbb{E}[\cdot]$ represents the expectation taken with respect to the filtration $\mathcal{F}_{k}$. It is important to note that the iterates $u_{k}^{\delta}$ are measurable with respect to $\mathcal{F}_{k}$ (see [15] for more details).
The following well-known identities will be frequently utilized: for any random variables $\Theta,\Theta_{1},\Theta_{2}$ and $\Phi$, and real constants $c_{1},c_{2}$, we have
  • (a)
    $\mathbb{E}[\Theta]=\mathbb{E}[\mathbb{E}[\Theta|\mathcal{F}_{k}]].$
  • (b)
    Additive: $\mathbb{E}[\Theta_{1}+\Theta_{2}|\mathcal{F}_{k}]=\mathbb{E}[\Theta_{1}|\mathcal{F}_{k}]+\mathbb{E}[\Theta_{2}|\mathcal{F}_{k}].$
    Linearity: $\mathbb{E}[c_{1}\Theta+c_{2}|\mathcal{F}_{k}]=c_{1}\mathbb{E}[\Theta|\mathcal{F}_{k}]+c_{2}.$
  • (c)
    If $\Theta\leq\Phi$, then $\mathbb{E}[\Theta|\mathcal{F}_{k}]\leq\mathbb{E}[\Phi|\mathcal{F}_{k}]$ almost surely.
Here the expression $\mathbb{E}[\Theta|\mathcal{F}_{k}]$ represents the expected value of $\Theta$ given the information contained in $\mathcal{F}_{k}$.
Triangular inequality for expectation: Let $\Theta_{1}$ and $\Theta_{2}$ be random variables. Then, there holds
((2.1)) $\mathbb{E}[\|\Theta_{1}+\Theta_{2}\|]\leq\mathbb{E}[\|\Theta_{1}\|]+\mathbb{E}[\|\Theta_{2}\|].$
Jensen’s inequality: For any random variable $\Theta$ and for any convex function $f$, there holds
((2.2)) $f(\mathbb{E}[\|\Theta\|])\leq\mathbb{E}[f(\|\Theta\|)],$
provided $\mathbb{E}[\|\Theta\|]<\infty$ and $\mathbb{E}[f(\|\Theta\|)]<\infty$ (see [1] for more details).
Next, we introduce an important result that splits the mean squared error $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]$, where $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$ represents the error for any iteration $u_{k}^{\delta}$. This decomposition is a fundamental strategy in error analysis and plays a crucial role in deriving the convergence rates for iterative methods.
Bias-variance decomposition: The bias-variance decomposition for the iterative scheme is given by
((2.3)) $\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]=\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}+\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}],$
where $B$ is defined as in (4.1) and $h\geq 0.$ In this decomposition:
  • The first term, $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}$, represents the bias, which includes both the approximation error and the error due to noise in the data.
  • The second term, $\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}]$, represents the variance, which accounts for the randomness introduced by the selection of the index $i_{k}$ during the iterations.
The bias reflects the deterministic error from the method’s approximation, while the variance captures the stochastic error due to the randomized nature of the iterative process.
We end this section by discussing few notations.
For some $u\in U$ and $\rho>0,$ we denote the open and closed balls in $U$ of radius $\rho$ centered at $u$, respectively by $B_{\rho}(u)$ and $\overline{B}_{\rho}(u).$ Let $\mathcal{D}_{\rho}(u^{\dagger})$ denote the set of all the solutions of (1.2) in $\overline{B}_{\rho}(u^{\dagger})$, i.e., $\mathcal{D}_{\rho}(u^{\dagger}):=\{u\in\overline{B}_{\rho}(u^{\dagger}):F(u)=y^{\dagger}\}$.

### v3

@@
Abstract. Over the past decade, stochastic algorithms have emerged as scalable and efficient tools for solving large-scale ill-posed inverse problems by randomly selecting subsets of equations at each iteration. However, due to the ill-posedness and measurement noise, these methods often suffer from oscillations and semi-convergence behavior, posing challenges in achieving stable and accurate reconstructions. This study proposes a novel variant of the stochastic gradient descent (SGD) approach, the iteratively regularized stochastic gradient descent (IRSGD) to address nonlinear ill-posed problems in Hilbert spaces. Under standard assumptions, we demonstrate that the mean square iteration error of the method tends to zero for exact data. In the presence of noisy data, we first propose a heuristic parameter choice rule (HPCR) and then apply the IRSGD method in combination with HPCR. Precisely, HPCR selects the regularization parameter without requiring any a-priori information of the noise level. We show that the method terminates in finitely many steps in case of noisy data and has regularizing features. Finally, some numerical experiments are performed to demonstrate the practical efficacy of the method.
Keywords: Unknown noise level; Stochastic gradient descent; Heuristic stopping rules; Nonlinear ill-posed problems; Inverse problems; Schlieren tomography.
MSC codes: 65J22, 65J20, 47J06
@@ 1. Introduction
This work focuses on deriving an approximate solution for a system of nonlinear ill-posed equations
((1.1)) $F_{i}(u)=y^{\dagger}_{i},$
where $i=0,1,\ldots,P-1,$ and for each $i$, the mapping $F_{i}:\mathfrak{D}(F_{i})\subset U\rightarrow Y_{i}$ represents a nonlinear operator between the Hilbert spaces $U$ and $Y_{i}$ with domain $\mathfrak{D}(F_{i})$. In this context, $y_{i}^{\dagger}\in Y_{i}$ represents the exact data, $P\in\mathbb{N}$, $U$ and $Y_{i}$ are endowed with their standard inner products and norms, respectively. Systems of the form (1.1) commonly appear in various real-life applications. For instance, they are fundamental in various tomography methods used in fields such as medical imaging and industrial diagnostics [7, 20, 29]. Conventionally, for the space $Y^{P}=Y_{0}\times Y_{1}\times\cdots\times Y_{P-1}$ equipped with the natural inner product derived from those of $Y_{i}$, (1.1) can be interpreted as
@@
where the mapping $F:U\rightarrow Y^{P}$ is given by
$F(u)=\begin{pmatrix}F_{0}(u)\\ \vdots\\ F_{P-1}(u)\end{pmatrix},\ \ \text{and}\ \hskip 5.69054pty^{\dagger}=\begin{pmatrix}y^{\dagger}_{0}\\ \vdots\\ y^{\dagger}_{P-1}\end{pmatrix}$.
@@
The exact data $y^{\dagger}$ is typically never available. Rather, a noisy data $y^{\delta}\in Y^{P}$ with the noise level $\delta\geq 0$ is accessible such that
((1.3)) $\|y^{\dagger}-y^{\delta}\|\leq\delta.$
Here, we write $\delta=\sqrt{\delta_{0}^{2}+\delta_{1}^{2}+\cdots+\delta_{P-1}^{2}}$ and $y^{\delta}=(y^{\delta}_{0},y^{\delta}_{1},...,y^{\delta}_{P-1})$, where
((1.4)) $\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\delta_{i}\ \ \text{for}\ \ i=0,1,\ldots,P-1.$
@@
The solution to (1.1) might not exist because of its ill-posedness, and even if it does, it is unlikely to be unique. Moreover, the solution(s) can be highly sensitive to perturbations in $y^{\delta}$ (see [5]). Therefore, regularization techniques are required to find the stable approximate solution of (1.1).
@@
Iterative regularization is a well-established and highly effective class of regularization techniques. This approach has achieved significant success in solving a wide variety of inverse problems, (see [5, 22]). Among these methods, the Landweber Iteration Method (LIM) is a widely recognized classical iterative technique. For the mapping $F:U\rightarrow Y^{P}$ that is Fréchet differentiable, LIM is
((1.5)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta}),\hskip 28.45274ptk\geq 0,$
@@
where $u_{0}^{\delta}:=u_{0}$ is an initial guess and $F^{\prime}(u_{k}^{\delta})^{*}$ represents the adjoint of the Fréchet derivative of $F(u_{k}^{\delta})$.
@@
The method (1.5) was modified by Scherzer [33], who included a damping term and called it the iteratively regularized Landweber iteration method. For $\lambda_{k}\geq 0$, the iteration takes the form
((1.6)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta})-\lambda_{k}(u_{k}^{\delta}-u^{(0)}),$
where $k\geq 0$. The $k$-th iteration of ((1.6)) can be viewed as performing a gradient descent step on the functional
$J(u)=\frac{1}{2}\left(\|F(u)-y^{\delta}\|^{2}+\lambda_{k}\|u-u^{(0)}\|^{2}\right)=\frac{1}{2}\left(\sum_{i=0}^{P-1}\|F_{i}(u)-y_{i}^{\delta}\|^{2}+\lambda_{k}\|u-u^{(0)}\|^{2}\right).$
Compared to the method (1.5), the analysis of deriving convergence rates for the method (1.6) requires less assumptions on $F$. Moreover, the method (1.6) converges to a solution that remains close to $u^{(0)}$, whereas achieving a similar convergence behavior with the method (1.5) needs additional assumptions (see [22]). This highlights the advantage of incorporating the extra damping term in (1.6) as compared to (1.5). When the noisy data is available, it is necessary to use an appropriate stopping rule (a-posteriori) to illustrate that the iterative schemes (1.5) and (1.6) are regularization methods. The prevalent stopping rule is referred to as discrepancy principle [28], which states that the method is stopped after $k_{\delta}=k_{\delta}(\delta,y^{\delta})$ steps, where
@@
for some $\tau>1.$ However, it is challenging to apply the discrepancy principle (1.7) in practicality since the noise level is not always reliable or available. Consequently, while applying the discrepancy principle, overestimation or underestimation of the noise level could lead to serious deterioration of the reconstruction accuracy. Therefore, it is reasonable to work on heuristic stopping rules that do not depend on the information of noise level. Real [32] proposed a heuristic rule for the Landweber iteration (1.5) based on the discrepancy principle, which is further motivated by the method of Hanke and Raus [9]. For (1.5), heuristic rule determines an integer $k_{*}:=k_{*}(y^{\delta})$ by minimizing
@@
over $k$, where $p>0$ is a constant and $a\geq 1$ is a fixed number.
It may be observed that the significant computational cost per iteration is a common problem with traditional iterative regularization techniques mainly due to the need to process complete dataset. Applying stochastic gradient descent (SGD) [15, 16, 17, 31] is a viable and successful way to deal with this problem. Using [16], the SGD method can be expressed as
@@
where $i_{k}$ is a uniformly drawn random index from $\{0,1,\ldots,P-1\}$ and $\mu_{k}\geq 0$ is the step size. For the SGD method (1.9), the stopping index $k(\delta)\in\mathbb{N}$ satisfies the rule
@@
Notably, the aforementioned rule is an a-priori rule. We refer to [16, 24] for some recent work on the SGD method using a-priori stopping rule and [13, 35] for the SGD method using a-posteriori stopping rule. To the best of our knowledge, convergence results of SGD method via heuristic stopping rules are not yet available in the literature. Accordingly, the primary objective of this study is to explore the convergence analysis of the following iteratively regularized stochastic gradient descent (IRSGD) under heuristic stopping rules:
@@
where $\lambda_{k}$ is the weighted parameter, $\omega_{k}$ is the step size. This method differs from the SGD method in the sense that an extra damping term $\lambda_{k}(u_{k}^{\delta}-u_{0})$ is considered in its formulation. A key advantage of studying IRSGD lies in its ability to reflect SGD by setting the damping factor $\lambda_{k}=0$. Our approach is novel in the realm of stochastic methods for solving ill-posed problems as it provides convergence analysis without relying on an a-priori stopping rule. Additionally, it circumvents the need to know the noise level, which further enhances its practical applicability.
We perform the detailed convergence analysis of the proposed method with early stopping, which is guaranteed to terminate after a finite number of iterations. Unlike the a priori stopping rules, which require a deterministic stopping index, our method yields a stopping index that is inherently random due to its dependence on the sample paths. This randomness introduces analytical challenges that cannot be addressed using the techniques in [16, 20, 24]. To overcome these, we develop a convergence framework grounded in Hilbert space theory and probability.
Our convergence framework, which addresses the random stopping index defined by a heuristic rule, is inspired by the recent work [11], where random stopping based on the discrepancy principle is studied in the context of stochastic mirror descent. By constructing a probability-one event, pathwise convergence is established for the exact data case, and, together with a stability result, this leads to almost sure convergence in the noisy case as the noise level approaches zero.
The main outcomes of this study are listed as follows:
  • The proposed method is fully data-driven, requiring neither a-priori nor a-posteriori stopping rules. To the best of our knowledge, this is the first work to formulate and analyze a stochastic method with a heuristic-type stopping criterion.
  • Unlike prior works [13, 15, 16, 17] that employ a polynomially decaying step size, we allow the step size $\omega_{k}$ to vary arbitrarily within a fixed interval $\omega,\Omega]$. We establish convergence of the proposed IRSGD scheme ((1.11)) in Hilbert spaces under a heuristic stopping rule. The analysis is technically challenging due to the randomness of the stopping index and the inclusion of a damping term.
  • We extend the scope from analyzing a single equation to developing and studying stochastic methods for solving large-scale systems of equations (i.e., with large $P$ in ((1.1))). By randomly selecting and processing individual equations from ((1.1)) independently, our approach drastically reduces memory requirements, making it particularly relevant to large-scale inverse problems arising in practical applications [7, 29, 30].
  • Through extensive numerical experiments, we demonstrate that incorporating a damping factor into the SGD framework along with the heuristic stopping significantly improves both reconstruction quality and convergence speed compared to standard SGD variants, confirming the practical benefits of the proposed approach.
Outline. Section 2 introduces the required assumptions, presents the IRSGD algorithm, and develops the analysis for the heuristic stopping rule, including a result ensuring the iterates remain within a certain ball. Section 3 provides convergence results for both noisy and exact data. Section 4 reports numerical experiments illustrating the effectiveness of the method, and the final section concludes the paper.